# Learning to Collaborate: A Capability Vectors-based Architecture for Adaptive Human-AI Decision Making

### v1

@@ Abstract
Human-AI collaborative decision making has emerged as a pivotal field in recent years. Existing methods treat human and AI as different entities when designing human-AI systems. However, as the decision capabilities of AI models becoming closer to human beings, it is necessary to build a uniform framework for capability modeling and ̵integrating. In this study, we propose a general architecture for human-AI collaborative decision making, wherein we employ learnable capability vectors to represent the decision-making capabilities of both human experts and AI models. These capability vectors are utilized to determine the decision weights of multiple decision makers, taking into account the contextual information of each decision task. Our proposed architecture accommodates scenarios involving multiple human-AI decision-makers with varying capabilities. Furthermore, we introduce a learning-free approach to establish a baseline using global collaborative weights. Experiments on image classification and Hatespeech detection demonstrate that our proposed architecture significantly outperform the current state-of-the-art methods in image classification task and sentiment analysis, especially for the case with large non-expertise capability level. Overall, our method provides an effective and robust collaborative decision-making approach that integrates diverse human/AI capabilities within a unified framework.
@@ 1 Introduction
With the fast increasing of AI capability in recent years, human-AI collaborative system has become an important topic of study. Especially in the era of large language models, AI model can do much more language-based tasks with surprisingly high quality (OpenAI, 2023), which makes human-AI collaboration a promising way in improving the productivity in a variety of subjects.
@@
Existing studies have shown that these collaborative system can provide better decision quality compared with human-only or AI-only systems. For example, in skin cancer recognition, it is found that good quality AI-based support of clinical decision-making improves diagnostic accuracy over that of either AI or human doctors alone, and that the least experienced clinicians gain the most from AI-based support (Tschandl et al., 2020). In general image recognition, human-AI collaboration can provide a performance improvement from both human-only or AI-only system through learning to deter (Madras et al., 2018; Mozannar and Sontag, 2020; Verma and Nalisnick, 2022; Dvijotham et al., 2023) or probability merging (Kerrigan et al., 2021; Singh et al., 2023). For more complex tasks such as text generation, empirical evidence has shown that human-AI collaboration can improve both the efficiency and quality of writing (Noy and Zhang, 2023).
However, most existing studies on collaborative decision-making focus only on a limited subset of scenarios and could fail to generalize across various decision-makers with differing capabilities. Therefore, we argue that it is essential to develop a unified architecture capable of accommodating a wide range of collaborative decision-making scenarios, including multiple decision-makers, output merging, and decision-maker selection. Traditionally, research on human-AI collaborative decision-making treats humans and AI models as independent entities. Yet, recent advancements in AI systems, particularly large language models (LLMs), have led to the acquisition of skill sets that closely resemble those of human decision-makers (OpenAI, 2023). As a result, the gap between the decision-making capabilities of humans and AI is narrowing. In fact, for some particular types of decision tasks, it is even smaller than the difference between different human beings. To uniformly model the decision related capabilities of different humans and AI models, we introduce capability vectors, which is comparable to user embedding in recommendation systems (Zhang et al., 2019), and can be either pre-defined or learnt. Based on capability vectors, we build a joint collaborative decision making architecture for selection-based decision tasks. This architecture can be applied for the scenario of multiple decision makers, and applies a weighted merging of output vectors from all decision makers involved. Experiments demonstrate the effectiveness of the proposed architecture and use of capability vectors. The main contributions of our study can be given as follows:
  • We introduce the learnable capability vectors to uniformly model the decision related capabilities of both different human experts and different AI models.
  • We develop a learning method that can train the capability vectors as well as generating the decision weights of a set of decision makers for generating the final decision score.
  • We propose a generalized architecture for hybrid human-AI decision making based on a weighted combination of scores from multiple decision makers, capable of handling the decision of classification or generation models.
  • Additionally, we introduce a learning-free baseline approach by converting model output into a one-hot vector plus random noises, and apply a probabilistic model for searching the optimal global collaborative weights and predicting the final accuracy.
@@ 2.1 Collaboration Mechanisms
Traditionally, many studies on human-AI collaboration focus on using a model to determine whether to apply AI or human experts to get the final decision. Madras et al. (2018) proposed a two-stage framework containing an automated model and an external decision-maker as in rejection learning (Cortes et al., 2016; Zhang et al., 2023), and generalizes rejection learning by considering the effect of other agents in the decision-making process. Mozannar and Sontag (2020) further explored the method to learn a predictor that can either predict or choose to defer the decision task to a downstream expert. They also developed a novel reduction to cost sensitive learning where a consistent surrogate loss is given. Verma and Nalisnick (2022) pointed out that the multi-class framework in (Mozannar and Sontag, 2020) is not calibrated with respect to expert correctness, and propose an L2D system based on one-vs-all classifiers that is able to produce calibrated probabilities of expert correctness. Although many studies focus on building calibrated predictors, Vodrahalli et al. (2022) show that even when the original AI model is well-calibrated, AI models as more confident than they actually are can improve human-AI performance. Moreover, Bansal et al. (2021) argue that the most accurate AI is not necessary the best teammate in AI advised decision making, and AI systems should be trained in a human-centered manner, directly optimized for team performance. Kerrigan et al. (2021) develop a set of methods that combine the probabilistic output of a model with the class-level output of a human. They show that the accuracy of the combination model is driven by not only the accuracies of individual human and model, but also by the model’s confidence.
Some other works are dedicated in the case of ddecision makers. Hemmer et al. (2022) point out that in real-world situations several human experts with varying capabilities may be available, and develop a method that trains a classifier to complement the capabilities of multiple human experts by allocating each instance to the most suitable team member. Singh et al. (2023) show that combining multiple human labels with the model’s probabilistic output can lead to significant improvement in accuracy, which is achieved by an approach to merge the predicted labels from multiple humans with the model’s probabilistic output. Da and Huang (2020) investigates how herding behavior affects the accuracy of consensus forecasts on a crowd-based corporate earnings platform, and concludes that encouraging independent forecasting and limiting information access can enhance the wisdom of crowds. He et al. (2022) conducts a comprehensive comparison of 58 prominent models of risky choice using 19 behavioral datasets, revealing that model crowds outperform individual models in predicting choice behavior, suggesting that combining multiple models provides a more accurate understanding of decision-making processes. Zou et al. (2024) introduces a statistical method for assessing ensembles of any size and investigates the optimal human-machine ratio as well as the impact of ensemble size on the performance of human-machine collaboration. Mao et al. (2024b) presents a study of surrogate losses and algorithms for the general problem of learning to defer with multiple experts. Mao et al. (2024a) further explores a two-stage learning scenario for deferring decisions to multiple experts, where a predictor is trained using a common loss function, and subsequently a deferral function is learned to assign inputs to the most suitable expert.
Some papers focus on theoretical analysis of collaborative decision making. Lamberson and Page (2012) analyzes the optimal composition of a forecasting group based on type coherence, where forecasters of the same type share expected accuracy and higher covariance. Steyvers et al. (2022) presents a Bayesian framework for integrating predictions and confidence scores from both humans and AI algorithms to enhance performance in classification tasks, which demonstrates that hybrid systems can outperform individual human or machine predictions even if their accuracy levels differ, as long as these differences are within a certain bound. Zou et al. (2024) explores the dynamics of multi-human and multi-machine interactions within hybrid human-machine systems and introduces a novel statistical framework to assess integration accuracy.
2.2 Capability Modeling.
@@
There exists many metrics for evaluating human capabilities. For instance, the Intelligence Quotient (IQ) and Emotional Intelligence Quotient (EQ) are widely utilized to quantify the two major aspects of human intelligence (Cattell, 1943; Salovey et al., 2001). In addition to these single-value metrics, various studies have employed multi-dimensional representations to model human personality traits, such as the Big Five personality model and the Myers-Briggs Type Indicator (Goldberg, 2013; Pittenger, 1993; Myers et al., 1999). Furthermore, language proficiency assessments like TOEFL, IELTS, and PETS, along with professional competency examinations such as the Bar Exam and CPA exam, evaluate human capabilities across different domains. Simultaneously, there is a growing body of research focused on assessing the capabilities of AI models. Task-specific models are typically evaluated using metrics relevant to their respective tasks, such as accuracy or recall for classification tasks, and BLEU or ROUGE scores for language generation tasks. For models designed to perform multiple tasks, their capabilities can be assessed through a combination of metrics derived from these various tasks. Additionally, large language models (LLMs), which are capable of executing a wide range of language-based tasks akin to human performance, often require human evaluation for comprehensive assessment. Numerous benchmarks have been established for evaluating LLMs (Chang et al., 2024). Some studies have even explored the application of IQ as a metric for evaluating AI models (Liu et al., 2019; Holzinger et al., 2019; Kim et al., 2021).
@@
As AI models increasingly exhibit capabilities that overlap with those of humans, it becomes imperative to develop unified evaluation metrics for both entities. Such metrics should extend beyond singular indices like IQ and may take the form of learnable vectors, akin to token embeddings used in natural language processing. We believe this is crucial for developing a robust human-AI collaborative system with enhanced generalization capabilities.
3 Preliminary
This study applies a transformer encoder to build the combination weights from decision makers. Transformer is a widely utilized deep learning architecture for processing sequential data (Vaswani, 2017), where the input for each sequential step is represented by an embedding vector. Let the input sequence be denoted as $\mathbf{X}=[\mathbf{x}_{1},...,\mathbf{x}_{n}]$, where each $\mathbf{x}_{i}$ is a vector. The dimension of the input sequence is $n\times d_{e}$, where $d_{e}$ is the dimension of each embedding vector $\mathbf{x}_{i}$ and $n$ is the sequence length. For language, visual or time series data, usually positional encoding is needed as model parameters of transformers do not distinguish the absolute or relative position of the sequence. In addition, segment embeddings can be added for distinguishing segments of the input sequences. Therefore, the final input representation is given by: $\mathbf{Z}=\mathbf{X}+PE+SE$. A transformer encoder consists of a stack of multiple transformer encoder blocks. Each block comprises a multi-head self-attention mechanism and a point-wise feed-forward neural network, with two residual connections and layer normalization applied following both the multi-head attention and feed-forward operations. Ultimately, the output of a transformer encoder remains a sequence of vectors that retains the same dimensionality as $\mathbf{Z}$. Typically, linear transformations can be applied to a single output embedding or a pooled output embedding for regression or classification tasks.
4 Method
Existing studies have shown that combining the human prediction with model probabilities can lead to higher accuracies than model or human alone, while the parameters of the combination method can be estimated effectively (Kerrigan et al., 2021). It is also noticed that a naive combination of humans with AI model can lead to poor accuracy, and selecting a subset of humans and combine their labels in an intelligent strategy is important (Singh et al., 2023). However, although Singh et al. (2023) consider the combination method by the expertise of human experts, they use confusion matrix to model the expertise, which is not adaptive to more complex decision tasks or classification task with a large number of categories. Moreover, the confusion matrix is superficial representation of human’s capability for recognizing images. For example, color blindness can not be represented by the confusion matrix among image categories but can affect the human’s capability for recognizing or classification tasks.
@@ 4 Method
To better model the underlying capabilities of human or AI that may determine their decision performance in a collaborative system, we introduce learnable capability vectors. They can be applied in determining the optimal collaborative decision mechanism or decision weights for different decision tasks, and can be trained or pre-defined. This is comparable to the user or item embedding in recommendation systems (Yu et al., 2016; Barkan and Koenigstein, 2016; Zhang et al., 2019; Ghasemi et al., 2021), or word/token embedding in natural language processing (Kenton and Toutanova, 2019). We assume each candidate decision maker has a capability vector, and all candidate decision makers can formulate a vocabulary. Overall, our method is build upon the following three assumptions:
  • There are many AI models or human with different capabilities, which can be represented by capability vectors with the same set of dimensions.
  • The optimal collaborative mechanism or decision weight for each decision maker is not fixed, but should be determined along with the other decision makers as well as the decision task.
  • The performance of collaborative decision making is determined by capability vectors of decision makers, the decision task and the collaborative decision mechanism.
@@ 4.1 Model architecture
We explore a generalized framework for human-AI decision-making. Given a decision task associated with a decision context, we can get $n$ decision choices, which can be pre-defined or generated by a human-AI system. Our objective is to identify the optimal choice among these options by leveraging a combination of human experts and AI models as decision-makers. In this study, we propose to use a weighted linear combination of scores from all the decision makers to evaluate the goodness of each decision choice, and use that to rank the choices and make the final decision selection. The main architecture is shown in Figure  1.
Figure 1: The diagram of the proposed architecture. The candidate decision options may be either pre-defined or generated by models or human experts. We employ a weighted aggregation of the scores assigned by each decision maker to each option in order to derive the final score for each choice. The weights assigned to each decision maker are determined by a weight generation model, which takes into account the capability vectors and the embedding of the decision task.
Assume that we have $m$ decision makers, whose capability vectors are ${\mathbf{c}_{1},...,\mathbf{c}_{m}}$. Then the weight generation model can give the weighting scores:
$w_{1},w_{2},...,w_{m}=f(\mathbf{c}_{1},...,\mathbf{c}_{m};\mathbf{x};\mathbf{\theta})$ (1)
where $f(.)$ is the weights generation model. In this study, we use a transformer encoder for this component, while $\theta$ is the set of model parameter, and $x$ is the embedding of the decision task information. $w_{1},...w_{m}$ are the decision weights for the decision makers. Usually, the size of a capability vector is different from that of the embedding of the input decision context. Therefore, we need to add an extra linear transformation layer on the capability vectors or the task embeddings for aligning the dimensions $w_{i}=\mathbf{h}_{N_{i}}\mathbf{W}_{hw}$, $i=\{1,...,m\}$, where $\mathbf{W}_{hw}\in\mathbb{R}^{d\times 1}$ and $\mathbf{h}_{N_{i}}\in\mathbb{R}^{1\times d}$. Meanwhile, if the number of the decision makers $m$ is not fixed, we need to add a separation embedding between the capability embeddings and the decision context embeddings before being processed by the multi-head attention module. In addition, we apply a learnable positional embedding as in BERT to distinguish the positional information of each embeddings (Kenton and Toutanova, 2019).
In the side of decision makers, assume that there are $n$ possible decision outcomes, and each of them will be assessed by all of the decision makers and be assigned with $m$ scores. Therefore, we can get a score matrix $\mathbf{S}\in\mathbb{R}^{m\times n}$. Given the generated weights of the decision makers, the final scores of the candidate decision outcomes can be given by:
@@
where $\mathbf{w}=\{w_{1},w_{2},...,w_{m}\}\in\mathbb{R}^{1\times m}$, and $\mathbf{s}_{f}\in\mathbb{R}^{1\times n}$ is the vector of final scores of all the decision choices. We use a Softmax function to normalize the output, and then the weight of each category is given by:
$a_{i}=\frac{e^{s_{f_{i}}}}{\sum_{i}^{m}e^{s_{f_{i}}}}$ (3)
where $i$ is the index of each category. Assume that the labels of the optimal decision among the candidate decision outcomes are given in the format of one-hot vector ${y^{(1)},...,y^{(N)}}$. Then we can use a cross-entropy loss for model training:
@@
where $N$ is the number of examples in the labeled data batch. In general, the proposed architecture ensures that: (1) We apply a weighted combination of the decisions scores or decision selections from multiple decision makers with different capabilities, no matter humans or AI models, which include the case of probability combination as discussed in (Kerrigan et al., 2021). (2) The decision weights depend on both the decision task and the capabilities of decision makers in a continuous space, which can be generalized to out-of-sample decision makers given the uniformly defined capability vectors. (3) Both the weight generation model and the capability vectors are learnable in an end-to-end manner given sufficient training data.
In decision tasks like image classification, it is possible that different human experts have different capabilities. For example, an expert that is familiar with botany is more capable in recognizing botany types while an expert that is familiar with animals is more capable in recognizing animal types. In our method, the capability vectors will involve these difference between experts. For the weight generation model, it only need to recognize if a image is about a botany or an animal, and give the suitable expert more decision weights based on the capability vectors, while it does not need to make fine-grained recognition for the image. This make it possible to use a less powerful weight model to achieve high-accuracy collaborative decision making with existing model and human experts.

### v2

# Learning to Collaborate: A Capability Vector-based Architecture for Adaptive Human-AI Decision Making

@@ Abstract
Effective human–AI collaboration hinges on the ability to dynamically integrate the complementary strengths of human experts and AI models across diverse decision contexts. Context-aware weighted combination of human and AI outputs is a promising technique, which involves the optimization of combination weights based on capabilities of decision agents on a given task. However, existing approaches treat humans and AI as isolated entities, lacking a unified representation to model the heterogeneous capabilities of multiple decision agents. To address this gap, we propose a novel capability-aware architecture that models both human and AI decision-makers using learnable capability vectors. These vectors encode task-relevant competencies in a shared latent space and are used by a transformer-based weight generation module to produce instance-specific aggregation weights. Our framework supports flexible integration of confidence scores or one-hot decisions from a variable number of agents. We further introduce a learning-free baseline using optimized global weights for human–AI collaboration. Extensive experiments on image classification and hate speech detection tasks demonstrate that our approach outperforms state-of-the-art methods under various collaboration settings with both simulated and real human labels. The results highlight the robustness, scalability, and superior accuracy of our method, underscoring its potential for real-world applications.
@@ 1 Introduction
With the rapid advancement of AI capabilities in recent years, human-AI collaborative systems have become an increasingly important topic of study. The emergence of large language models (LLMs) has enabled AI to perform a wide range of language-based tasks with remarkably high quality (OpenAI 2023). However, AI systems still face limitations in areas requiring human-like nuanced judgment, real-world contextual understanding, value-based ethical considerations, or handling unstructured novel situations. Consequently, effectively integrating complementary human strengths with powerful AI capabilities presents a promising approach to enhance productivity and decision-making quality across diverse domains.
@@
Existing studies have shown that these collaborative system can provide better decision quality compared with human-only or AI-only systems. For example, in skin cancer recognition, it is found that good quality AI-based support of clinical decision-making improves diagnostic accuracy over that of either AI or human doctors alone, and that the least experienced clinicians gain the most from AI-based support (Tschandl et al. 2020). In general image recognition, human-AI collaboration improves performance over human-only or AI-only systems through methods such as learning to defer (Madras et al. 2018, Mozannar and Sontag 2020, Verma and Nalisnick 2022, Dvijotham et al. 2023) and probability merging (Kerrigan et al. 2021, Singh et al. 2023). For more complex tasks such as text generation, empirical evidence has shown that human-AI collaboration can improve both the efficiency and quality of writing (Noy and Zhang 2023).
However, most existing studies on collaborative decision-making focus on limited scenarios and could fail to generalize across various decision agents with different capabilities. Meanwhile, traditional research often treats humans and AI models as independent entities in collaborative decision-making. Yet, advances in AI systems, such as modern computer vision models and large language models (LLMs), have endowed them with capabilities increasingly comparable to human decision agents (He et al. 2015, Redmon et al. 2016, Han et al. 2022, OpenAI 2023, Luo et al. 2025), narrowing the gap between human and AI decision-making capacities. In addition, the diversity between human decision makers is also an issue to be concerned (Lu et al. 2024). To uniformly model human and AI agent decision-making within a multi-agent framework, we introduce capability vectors. Analogous to recommendation system user/item embeddings (Zhang et al. 2019), these vectors can be learned in an end-to-end manner. Building upon them, we develop a unified collaborative decision-making architecture supporting multiple human/AI agents via weighted merging of their output vectors. Experiments validate the effectiveness of both the architecture and the capability vectors.
The main contributions of our study include the following aspects: First, we introduce the learnable capability vectors to uniformly model the decision related capabilities of both different human experts and different AI models. In addition, we develop a learning method that can train the capability vectors as well as generating the decision weights of a set of decision agents for generating the final decision score. Second, we propose a unified architecture for hybrid human-AI decision-making. This architecture leverages a weighted combination of output vectors from multiple decision agents and is capable of handling any multi-agent collaborative classification scenario, accommodating both one-hot and confidence-based outputs. Third, we develop a learning-free baseline approach by transforming model outputs into one-hot vectors augmented with random noise. We then apply a probabilistic model to search for optimal global collaborative weights and predict the final accuracy.
The remainder of the paper is structured as follows. Section 2 reviews related work on human-AI collaboration mechanisms and capability modeling. Section 3 introduces our proposed methodology, including the architecture for weighted decision fusion, the learning mechanism for capability vectors, and an analysis of computational complexity. In section 4, we presents a learning-free baseline for multi-class classification using optimized global weights. Section 5 details extensive experiments on image classification and hate speech detection tasks, comparing our approach with state-of-the-art methods under various collaborative settings. In section 6, we provide an in-depth analysis of the model’s behavior, including the effect of non-expertise capability levels and the learning dynamics of capability vectors. We further discuss the broader implications of our work and potential application scenarios in Section 7. Finally, we concludes the paper and suggests directions for future research in Section 8.
@@ 2.1 Collaboration Mechanisms
Early work addressed AI deferring to single human experts, including two-stage rejection frameworks (Madras et al. 2018), consistent surrogate losses for deferral (Mozannar and Sontag 2020), and calibrated multiclass deferral (Verma and Nalisnick 2022). Key findings show that well-calibrated AI confidence enhances collaboration (Vodrahalli et al. 2022), optimizing individual AI accuracy does not necessarily optimize team performance (Bansal et al. 2021), and deferral notifications impact human-AI accuracy (Bondi et al. 2022). Complementary work includes taxonomies for analyzing complementary strengths (Rastogi et al. 2023), principled human teaching for deferral (Mozannar et al. 2022), and trust-driven reliance modeling (Li et al. 2023).
Recent advances extend collaboration to multiple decision agents. Kerrigan et al. (2021) integrated model probabilities with human class-level outputs, revealing dependencies on individual accuracies and confidence. Hemmer et al. (2022) pioneered team formation strategies for human experts with instance allocation mechanisms. Ensemble techniques were further explored through human-AI label integration (Singh et al. 2023), crowd wisdom principles (Da and Huang 2020), and model crowds (He et al. 2022). Methodological contributions include statistical ensemble assessment (Zou et al. 2024) and multi-expert deferral algorithms (Mao et al. 2024b, a).
2.2 Capability Modeling
@@
There exists many metrics for evaluating human capabilities. For instance, the Intelligence Quotient (IQ) and Emotional Intelligence Quotient (EQ) are widely utilized to quantify the two major aspects of human intelligence (Cattell 1943, Salovey et al. 2001). Beyond unidimensional metrics, studies frequently adopt multidimensional frameworks like the Big Five personality model and the Myers-Briggs Type Indicator to represent personality, a key factor shaping decision-making styles (Goldberg 2013, Pittenger 1993, Myers et al. 1999). Furthermore, language proficiency assessments like TOEFL, IELTS, and PETS, along with professional competency examinations such as the Bar Exam and CPA exam, evaluate human capabilities across different domains. Simultaneously, there is a growing body of research focused on assessing the capabilities of AI models. Task-specific models are typically evaluated using metrics relevant to their respective tasks, such as accuracy or recall for classification tasks, and BLEU or ROUGE scores for language generation tasks. For models designed to perform multiple tasks, their capabilities can be assessed through a combination of metrics derived from these various tasks. Additionally, large language models (LLMs), which are capable of executing a wide range of language-based tasks akin to human performance, often require human evaluation for comprehensive assessment. Numerous benchmarks have been established for evaluating LLMs (Chang et al. 2024). Some studies have even explored the application of IQ as a metric for evaluating AI models (Liu et al. 2019, Holzinger et al. 2019, Kim et al. 2021).
@@
As AI models increasingly exhibit capabilities that overlap with those of humans, it is of great value to build unified evaluation metrics for both entities. Such metrics should extend beyond singular indices like IQ and may take the form of learnable vectors, akin to token embeddings used in natural language processing, or user embeddings in recommender systems.
3 Methodology
Human-AI collaborative decision-making has two primary objectives: enhancing the accuracy and quality of decisions, and reducing the associated decision-making costs. This study focuses on the former, specifically aiming to improve decision accuracy in classification tasks. To this end, it is essential to incorporate not only the final decision choices but also the confidence scores across candidate labels. In practice, however, human decision makers typically provide only their selected choices, whereas AI models can output predicted probabilities for all possible options. When these probabilities are well calibrated, they can serve as reliable confidence scores.
Existing learning-to-defer (L2D) methods generally assign decisions either to a human or an AI model, without effectively integrating their respective confidence scores. Moreover, the cost associated with human involvement remains dependent on the deferral rate. Recent research has demonstrated that combining human predictions with model probabilities can yield higher accuracy than relying on either humans or the model alone, with parameters of such a combination being efficiently estimable (Kerrigan et al. 2021). At the same time, it has been observed that a naive integration of human and AI judgments may lead to suboptimal performance, underscoring the importance of selectively aggregating human inputs through intelligent strategies (Singh et al. 2023). However, while existing methods often combine decisions based on human expertise modeled via confusion matrices, such representations exhibit notable limitations. They lack adaptability to complex multi-agent decision environments or classification tasks with large label spaces. Furthermore, confusion matrices offer a purely empirical and often superficial mapping of expert behavior, failing to capture underlying cognitive patterns. For instance, conditions such as color blindness, which significantly impact human performance in visual classification, cannot be adequately represented in a confusion matrix, yet they critically influence decision-making ability.
To avoid the shortcomings of existing methods, we apply instance-aware weighted combination of confidence scores and one-hot decision choices over decision options. The combination weights is determined by the capabilities of decision agents on each task instance. To better model the underlying capabilities of human or AI that may determine their decision performance in a collaborative system, we introduce learnable capability vectors with the following properties:
  • A capability vector $\mathbf{C}\in\mathbb{R}^{1\times d}$ is a 1-D vector with pre-defined length $d$.
  • The capability vector reflects all of a decision agent’s capabilities related to a decision-making task category. It indicates that the subject’s expected accuracy in a specific task within the category can be entirely determined by the capability vector and the task’s attributes.
[…]
@@ 3 Methodology
They can be applied in determining the optimal collaborative decision mechanism or decision weights for different decision tasks, and can be trained or pre-defined. This is similar to the user or item embedding in recommendation systems (Yu et al. 2016, Barkan and Koenigstein 2016, Zhang et al. 2019, Ghasemi et al. 2021), or word/token embedding in natural language processing (Kenton and Toutanova 2019). We assume that each candidate agent is associated with a capability vector, and the set of all these capability vectors forms a vocabulary. Overall, our method is build upon the following three assumptions: (a) There are many AI models or human experts with different capabilities, which can be represented by capability vectors with the same set of dimensions. (b) The optimal collaborative mechanism or decision weights for each agent is not fixed, but should be determined along with the other agents as well as the decision task. (c) The performance of collaborative decision making is determined by capability vectors of decision agents, the decision task and the collaborative decision mechanism.
Figure 1: The diagram of the proposed architecture. The candidate decision options may be either pre-defined or generated by models or human experts. We employ a weighted aggregation of the scores assigned by each decision agent to each option in order to derive the final score for each choice. The weights assigned to each agent are determined by a weight generation model, which takes into account the capability vectors and the embedding of the decision task.
@@ 3.1 Model architecture
We explore a generalized framework for human-AI decision-making. Given a decision task associated with a decision context, we can get $n$ decision choices, which can be pre-defined or generated by a human-AI multi-agent system. Our objective is to identify the optimal choice among these options by leveraging a combination of human experts and AI models as decision agents. In this study, we propose to use a weighted linear combination of scores from all the decision agents to evaluate the goodness of each decision choice, and use that to rank the choices and make the final decision. The main architecture is shown in Figure  1.
Assume that we have $m$ decision agents, each having a capability vector denoted as $\mathbf{c}_{1},...,\mathbf{c}_{m}$}. Then the weight generation model can give the weighting scores:
$w_{1},w_{2},...,w_{m}$ $=f(\mathbf{c}_{1},...,\mathbf{c}_{m};\mathbf{x};\mathbf{\theta})$
$\sum_{i}^{m}w_{i}$ $=1$ (1)
where $f(.)$ is the weights generation model. Here we use a transformer encoder for this component (Vaswani 2017), while $\theta$ is the set of model parameter, and $x$ is the embedding of the decision task information. $w_{1},...w_{m}$ are the decision weights for the decision agents. Usually, the size of a capability vector is different from that of the embedding of the input context. Thus, we add an extra linear transformation layer on the capability vectors or the task embeddings for aligning the dimensions $w_{i}=\mathbf{h}_{N_{i}}\mathbf{W}_{hw}$, $i=\{1,...,m\}$, where $\mathbf{W}_{hw}\in\mathbb{R}^{d\times 1}$ and $\mathbf{h}_{N_{i}}\in\mathbb{R}^{1\times d}$. Meanwhile, if the number of the decision agents $m$ is not fixed, a separation embedding between the capability embeddings and the decision context embeddings is needed. In addition, we apply a learnable positional embedding as in BERT to distinguish the positional information of each embeddings (Kenton and Toutanova 2019).
In the side of decision agents, assume that there are $n$ possible decision outcomes, and each of them will be assessed by all of the agents and be assigned with $m$ scores. Therefore, we can get a score matrix $\mathbf{S}\in\mathbb{R}^{m\times n}$. Given the generated weights of the decision agents, the final scores of the candidate decision outcomes are given by:
@@
where $\mathbf{w}=\{w_{1},w_{2},...,w_{m}\}\in\mathbb{R}^{1\times m}$, and $\mathbf{s}_{f}\in\mathbb{R}^{1\times n}$ is the vector of final scores of all the decision choices. We use the Softmax function to normalize the output, and then the weight of each category is given by $a_{i}=\frac{e^{s_{f_{i}}}}{\sum_{i}^{m}e^{s_{f_{i}}}}$, where $i$ is the index of each category. Assume that the labels of the optimal decision among the candidate decision outcomes are given in the format of one-hot vector $y^{(1)},...,y^{(N)}$}. Then we can use a cross-entropy loss for model training:
@@
where $N$ is the number of examples in the labeled data batch. In general, the proposed architecture ensures that: (1) We apply a weighted combination of the decisions scores or decision selections from multiple decision agents with different capabilities, no matter humans or AI models, which include the case of probability combination as discussed in (Kerrigan et al. 2021). (2) The decision weights depend on both the decision task and the capabilities of decision agents in a continuous space, which can be generalized to out-of-sample decision agents given the uniformly defined capability vectors. (3) Both the weight generation model and the capability vectors are learnable in an end-to-end manner given sufficient training data.